OpenStudy (anonymous):

Circle A is shown with a central angle marked 37 degrees and the radius marked 10 inches. Which of the following could be used to calculate the area of the sector in the circle shown above? π(10in)237 over 360 π(10in)37 over 360 π(37in)210 over 360 π(37in)10 over 360

OpenStudy (anonymous):

im a noob a math

OpenStudy (anonymous):

*at*

jimthompson5910 (jim_thompson5910):

Hint: area of slice = \(\LARGE \frac{x}{360}*\pi*r^2\) x is the angle in degrees r is the radius

OpenStudy (anonymous):

i don't understand how to use pi in equations

jimthompson5910 (jim_thompson5910):

what I would do is pull out the pi and ignore it temporarily

jimthompson5910 (jim_thompson5910):

In this case, x = 37 degrees and r = 10 if r = 10, then what is the value of r^2 ?

OpenStudy (anonymous):

100

OpenStudy (anonymous):

so x*r^2 is 3700?

OpenStudy (anonymous):

someone help me

OpenStudy (anonymous):

:(

jimthompson5910 (jim_thompson5910):

so, \[\Large \frac{x}{360}*\pi*r^2 = \frac{37}{360}*\pi*10^2\] \[\Large \frac{x}{360}*\pi*r^2 = \frac{37}{360}*\pi*100\] then you can rearrange terms

jimthompson5910 (jim_thompson5910):

hmm strange how none of the answer choices match

OpenStudy (anonymous):

well i don't know which one

OpenStudy (anonymous):

i asked one of my friends and he has no clue

jimthompson5910 (jim_thompson5910):

can you post a screenshot of the full problem?

OpenStudy (anonymous):

how im using a mac i don't know how

jimthompson5910 (jim_thompson5910):

http://www.take-a-screenshot.org/ click on the "mac" tab to have the instructions shown

OpenStudy (anonymous):

here?

OpenStudy (anonymous):

wait

jimthompson5910 (jim_thompson5910):

on the link I posted. They show the instructions on how to take a screenshot

OpenStudy (anonymous):

it doesn't work i tried all of them

jimthompson5910 (jim_thompson5910):

did you hold the command key and shift key, then press the "3" key? (the '3' over the 'w' and 'e') you should see a new pic file show up on your desktop

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

ok post that by attaching the file (click the attach file button under the text box)

OpenStudy (anonymous):

well no picture but i tried and it didn't work

jimthompson5910 (jim_thompson5910):

the file name will be something like `Screen shot 2015-08-31 at 08.45.00 AM.png`

OpenStudy (anonymous):

no its not working

jimthompson5910 (jim_thompson5910):

ok nvm then

OpenStudy (anonymous):

what do i do now with the equation?

jimthompson5910 (jim_thompson5910):

sorry I don't know. None of the answer choices match what I wrote above, so I'm stumped. I think there's a typo maybe?

jimthompson5910 (jim_thompson5910):

that's why I wanted to see the screenshot so I could see if anything was missing or not

OpenStudy (anonymous):

yay

jimthompson5910 (jim_thompson5910):

can you post an image of the answer choices as well?

OpenStudy (anonymous):

only like that π(10in)237 over 360 π(10in)37 over 360 π(37in)210 over 360 π(37in)10 over 360

OpenStudy (anonymous):

the question says: Which of the following could be used to calculate the area of the sector in the circle shown above?

jimthompson5910 (jim_thompson5910):

does any of the answer choices have \(\Large 10^2\) in it?

OpenStudy (anonymous):

the first 2

jimthompson5910 (jim_thompson5910):

so choice B is \[\Large \frac{\pi(10^2)37}{360}\] ??

OpenStudy (anonymous):

no lemme make that equation wait a sec

OpenStudy (anonymous):

crud its not working its just pi*10*(37/360)

OpenStudy (anonymous):

or the 2nd one just with 10^2 instead of 10

jimthompson5910 (jim_thompson5910):

ok, so the "10" should be "10^2" because of the r^2 in pi*r^2

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

look: π(10in^2)237 over 360 π(10in)37 over 360 π(37in^2)210 over 360 π(37in)10 over 360

OpenStudy (anonymous):

are you here?

jimthompson5910 (jim_thompson5910):

so strange, it shouldn't be 10in^2 it should be 10^2 in^2 maybe that's what they meant?